OpenStudy (anonymous):

Bill empties his jar of coins onto the floor. If the probability of all the coins landing on tails is 1/64, then what is the best answer for how many coins Bill has?

OpenStudy (anonymous):

Each coin has a 1/2 chance to land on tails. The probability for all of them landing on tails is: \[P_1(tails) \times P_2(tails) \times ... \times P_n(tails)\]\[=\frac{1}{2} \times\frac{1}{2} \times ... \times \frac{1}{2} = (\frac{1}{2})^n\] Where n is the number of coins. So if the probability is \(\frac{1}{64}\) then we have that \[(\frac{1}{2})^n = \frac{1}{64}\] So what is n?

OpenStudy (anonymous):

i dont know what is n

OpenStudy (anonymous):

Did you try it yet?

OpenStudy (anonymous):

yes i got it

OpenStudy (anonymous):

What did you get?